# Group-aware Parameter-efficient Updating for Content-Adaptive Neural Video Compression

### v1

@@ Abstract.
Content-adaptive compression is crucial for enhancing the adaptability of the pre-trained neural codec for various contents. Although these methods have been very practical in neural image compression (NIC), their application in neural video compression (NVC) is still limited due to two main aspects: 1), video compression relies heavily on temporal redundancy, therefore updating just one or a few frames can lead to significant errors accumulating over time; 2), NVC frameworks are generally more complex, with many large components that are not easy to update quickly during encoding. To address the previously mentioned challenges, we have developed a content-adaptive NVC technique called Group-aware Parameter-Efficient Updating (GPU). Initially, to minimize error accumulation, we adopt a group-aware approach for updating encoder parameters. This involves adopting a patch-based Group of Pictures (GoP) training strategy to segment a video into patch-based GoPs, which will be updated to facilitate a globally optimized domain-transferable solution. Subsequently, we introduce a parameter-efficient delta-tuning strategy, which is achieved by integrating several light-weight adapters into each coding component of the encoding process by both serial and parallel configuration. Such architecture-agnostic modules stimulate the components with large parameters, thereby reducing both the update cost and the encoding time. We incorporate our GPU into the latest NVC framework and conduct comprehensive experiments, whose results showcase outstanding video compression efficiency across four video benchmarks and adaptability of one medical image benchmark.
Neural Video Compression, Content Adaptive Neural Compression, Parameter-Efficient Transfer Learning
Corresponding Author: Dong Xu, dongxu@hku.hk
ccs: Computing methodologies Reconstructionccs: Computer vision problems
@@ 1. Introduction
Video compression, a cornerstone of multimedia computing, has been an essential technology for decades. This importance is driven by the need to efficiently store and transmit an ever-growing variety of video content. However, conventional video compression standards (e.g., H.266/VVC (Bross et al., 2021)) rely on manually designed modules that may no longer suffice for contemporary requirements. In recent times, Neural Video Compression (NVC) techniques (Li et al., 2023) that are optimized end-to-end with advanced neural modules have reached performance levels comparable to those of traditional codecs.
@@
Despite achieving significant success, there are still some limitations to NVC in practical coding applications. The primary challenge is generalization, as most existing NVC models struggle to perform adaptively across different sequential contents. This issue arises because most pre-trained NVC models are optimized for a specific expected rate-distortion (RD) cost within a certain domain (e.g., online video resource (Xue et al., 2019)), leading to a notable domain gap when applied to cases outside the training domain (e.g., medical volumetric images). To mitigate the issue of such domain gaps, several content-adaptive methodologies (Pan et al., 2022; Yang et al., 2020a; Campos et al., 2019; Xu et al., 2023; Zhao et al., 2021; Gao et al., 2022; Van Rozendaal et al., 2021; Jourairi et al., 2022; Lam et al., 2020; Zou et al., 2021; Tsubota et al., 2023; Shen et al., 2023; Wang et al., 2022) have been introduced in the field of neural image compression (NIC). Many approaches involve updating the parameters of the encoder neural modules (i.e., online-updating strategy) to generate quantized compressed features with improved rate-distortion values. This enhancement significantly improves compression performance in testing image scenarios.
@@
Conversely, adopting the direct migration of update strategies from NIC to NVC frameworks still seems impractical. A primary obstacle lies in the evolution of NVC architectures, which now incorporate extensive and complex modules (Li et al., 2021, 2022, 2023; Chen et al., 2023) with millions of parameters on the encoding side. Some early studies (Lu et al., 2020a) have explored online-updating all parameters on the encoding side mirroring strategies used in content-adaptive NICs. Though it initially yielded performance gains in early light-weight NVC frameworks (Lu et al., 2019), applying such updates to the intricate architectures of contemporary NVCs will result in prohibitively long encoding time and immense computational demands. Recent investigations (Tang et al., 2023) have also explored updating solely the motion to circumvent the necessity of optimizing all encoding-side parameters. However, given that motion usually represents just a minor portion of the total bit-rates, confining updates to merely motion components substantially limits the scope for performance enhancements.
@@
Another challenge limiting the effectiveness of content-adaptive NVC is error accumulation. Video compression heavily relies on temporal redundancy. i.e., the compression efficiency of the current frame is greatly dependent on the reconstruction quality of the previously compressed frame. However, the online updating strategy, which aims to optimize the rate-distortion value by minimizing both bit-rate and reconstruction quality simultaneously, can negatively impact the quality of subsequent coding frames through increased error accumulation. This dynamic underlines the difficulty of achieving both high compression efficiency and maintaining quality across frames.
@@
To tackle the aforementioned challenges, we propose a novel content-adaptive NVC method, termed Group-aware Parameter-efficient Updating (GPU). First, leveraging the breakthroughs in Parameter-Efficient Transfer Learning (PETL) (Ding et al., 2023; He et al., 2021; Houlsby et al., 2019) across various vision and language tasks, which demonstrates the efficiency of adapting a pre-trained model to new domains by fine-tuning a minimal set of parameters, our approach incorporates a low-rank adaptation style (Hu et al., 2021b) module, namely encoding-side adapters. These adapters are ingeniously crafted to stimulate specific components within the comprehensive encoding modules. By focusing optimization efforts on these adapters, which constitute a significantly smaller portion of the encoding modules’ parameters, we can then streamline optimization and reduce costs dramatically. Moreover, we delve into innovative ways to integrate these adapters into the NVC framework, exploring both serial and parallel configurations. The serial adapter is strategically placed immediately after a module, enhancing its processing, while the parallel adapter is woven into the module through a skip-connection layout, offering a blend of original and adapted processing pathways. This strategic placement of adapters allows for the modification of less than 10% of the parameters in our encoding module, markedly boosting compression efficiency with minimal modification.
@@
Second, to mitigate error accumulation, we introduce a group-aware optimization strategy. This method diverges from previous methods (Abdoli et al., 2023; Lu et al., 2020a; Tang et al., 2023), which concentrated on updating a narrow sequence of consecutive frames. Instead, our approach considers the entire Group of Pictures (GoP) during the updating process. Specifically, when compressing an individual frame, we optimize the entire GoP, which it belongs, to minimize the errors in individual reconstructed frames. Furthermore, to circumvent the significant memory demand that group-based updating could impose on the processing of high-resolution videos, we implement a patch-based GoP updating strategy. It spatially segments each GoP into multiple patch-based GoPs, which will be updated sequentially. In general, this approach ensures that the updates made to this framework do not adversely impact the predictive coding process of other frames within the GoP, thus maintaining the integrity and efficiency of the compression across the entire group.
@@
By integrating two above strategies, we incorporate our GPU into the latest NVC framework (Li et al., 2023), significantly improving its coding performance across four most representative video compression benchmarks, HEVC (Sullivan et al., 2012) Class B, C, D and E datasets. Moreover, to validate the adaptability of our content-adaptive method in various contents, we also test it on compressing medical volumetric image benchmarks ACDC (Bernard et al., 2018), which serves the most important cardiac dataset with multiple MRI slices. Our framework not only outperforms traditional video codecs, such as H.266/VVC (Bross et al., 2021), but also surpasses previous content-adaptive NVC methods (Lu et al., 2020a; Tang et al., 2023) in performance on video benchmarks and adaptability on medical benchmark. Our contributions can be summarized as follows:
@@
  • Implementation of a group-aware optimization strategy to mitigate error accumulation by considering an entire Group of Pictures (GoP) during online updating, ensuring consistent and improved predictive coding across frames.
  • Integration of our approach to enhance the coding performance of the latest NVC methods, demonstrating superior results over both traditional video codecs and existing content adaptive NVC methods across four key video and one medical volumetric image compression benchmarks.
@@ 2.1. Neural Video Compression
Current methods in neural video compression can be intuitively categorized into two main types: deep residual coding and deep contextual coding. The first-category approach is inspired by established video compression standards (Sullivan et al., 2012; Bross et al., 2021), which involves predictive coding (e.g., motion compensation) and the encoding of residual data. A notable advancement in this category is DVC (Lu et al., 2019), which utilizes Convolutional Neural Networks (CNNs) throughout the traditional residual coding process. Following this, numerous works (Hu et al., 2020, 2021a; Chen et al., 2022b; Lin et al., 2020; Yang et al., 2020b; Agustsson et al., 2020) have enhanced this framework by incorporating more robust modules and sophisticated approaches. For instance, FVC (Hu et al., 2021a) implements all coding operations in the feature space to improve compression efficiency.
@@
In the deep contextual coding category (Lombardo et al., 2019; Habibian et al., 2019; Chen et al., 2022a; Sheng et al., 2022; Li et al., 2022, 2023, 2024; Chen et al., 2023; Mentzer et al., 2022), research extends generative-based neural image compression (NIC) techniques by developing spatio-temporal conditional entropy models that leverage both spatial and temporal contexts. In the early studies, some works (Lombardo et al., 2019; Habibian et al., 2019) have introduced spatial-temporal conditional models by maintaining a latent temporal variable produced from all previous compressed frames. Unlike earlier approaches that rely on accumulating all available information, DCVC family (Li et al., 2021; Sheng et al., 2022; Li et al., 2022, 2023) introduces a series of methods that employ motion compensation from the consecutive compressed frame to derive temporal context directly, which has been further refined by progressive introducing new methods enhance such as multi-scale temporal context mining (Sheng et al., 2022) or hybrid entropy model (Li et al., 2022). The latest iteration of DCVC_DC has shown to surpass the performance of the conventional video coding standard H.266/VVC (Bross et al., 2021).
@@
Though above methods have achieved notable success, their performances are often constrained across diverse video content. This limitation stems primarily from the fact that many NVC frameworks are pre-trained on specific domains (Xue et al., 2019), which restricts their generalizability and adaptability to a wide range of content. To address this challenge, this work introduces a content-adaptive NVC method named GPU, which is seamlessly integrated into the latest NVC framework, DCVC_DC. This integration significantly enhances its compression performance, demonstrating the potential of content-adaptive approaches in overcoming the generalizability limitations of existing NVC frameworks.
2.2. Content Adaptive Neural Compression
@@
Content-adaptive methods play essential roles in both NIC (Pan et al., 2022; Yang et al., 2020a; Campos et al., 2019; Xu et al., 2023; Zhao et al., 2021; Gao et al., 2022; Van Rozendaal et al., 2021; Jourairi et al., 2022; Lam et al., 2020; Zou et al., 2021; Tsubota et al., 2023; Shen et al., 2023; Wang et al., 2022; Liu et al., 2023) and NVC (Lu et al., 2020a; Tang et al., 2023; Abdoli et al., 2023), which is to refine the latent representation produced by encoder side during encoding process. In this way, the model can adapt to current coding samples and reduce the domain shift between training and testing domains, which enhances the adaptability when compressing various types of testing data.
@@
Those approaches have yielded significant accomplishments in NIC. For instance, Campos et al. (Campos et al., 2019) refined latent variables through back-propagation, a technique that was further enhanced by Yang et al. (Yang et al., 2020a) through the incorporation of a differentiable quantization module. Concurrently, some researchers have pursued decoder-based strategies; for example, Rozendaal et al. (Van Rozendaal et al., 2021) opted to update all parameters within the decoder and the entropy model. Nonetheless, it’s important to note that these decoder-focused methods typically demand higher bit-rates for the storage and transmission of parameters to the decoder.
@@
In the field of NVC, the development of content-adaptive strategies is still in its infancy. Lu et al. (Lu et al., 2020b) introduced the first method of this kind by updating all parameters on the encoding side, drawing parallels with most content-adaptive approaches in NIC. This method was seamlessly integrated into early NVC framework (Lu et al., 2019), which typically featured lightweight architectures. However, as more advanced modules(e.g., Transformers (Vaswani et al., 2017)) are incorporated to enhance the performance, modern NVC frameworks (Li et al., 2021, 2022, 2023, 2024; Chen et al., 2023; Hu et al., 2022) have become significantly more complex. Updating the entire encoder is now impractical due to the substantial computational load and increased encoding time. To address this issue, Tang et al. (Tang et al., 2023) developed a novel methodology for online and offline updating, specifically targeting the compression of the motion. This approach requires motion information from H.266/VVC (Bross et al., 2021) as the auxiliary ground-truth. Yet, since motion data represents only a small fraction of the overall bit-rates, optimizing motion alone is insufficient for significant performance gains.
@@
In contrast, our method employs a group-aware and parameter-efficient updating strategy to overcome the aforementioned challenges. First, we jointly optimize the entire GoP when generating the bitstream for a single frame, which ensures that updating a particular frame does not lead to significant error accumulation. Secondly, we implement an encoder-side adaptor, which allows us to avoid optimizing the entire encoder network with those comprehensive modules. This adaptation makes the updating process more efficient in terms of parameter use.

### v2

@@ Abstract.
Content-adaptive compression is crucial for enhancing the adaptability of the pre-trained neural codec for various contents. Though, its application in neural video compression (NVC) is still limited due to two main aspects: 1), video compression relies heavily on temporal redundancy, therefore updating just one or a few frames can lead to significant errors accumulating over time; 2), NVC frameworks are generally complex, with many comprehensive components that are not trivial to update quickly during the encoding procedure. To address these challenges, we have developed a content-adaptive NVC technique called Group-aware Parameter-efficient Updating (GPU). Initially, to minimize error accumulation, we adopt a group-aware approach for updating encoder parameters. This involves adopting a patch-based Group of Pictures (GoP) updating strategy to segment a video into patch-based GoPs, which will be updated to facilitate a globally optimized domain-transferable solution. Subsequently, we introduce a parameter-efficient delta-tuning strategy, which is achieved by integrating several light-weight adapters into each encoding component by using both serial and parallel configuration. Such architecture-agnostic modules stimulate the components with large parameters, thereby reducing the updating cost during the encoding stage. We incorporate our GPU into the latest NVC framework and conduct extensive experiments, whose results showcase outstanding video compression efficiency across six video compression benchmarks and the adaptability of one medical volumetric image compression benchmark.
Neural Video Compression, Parameter-Efficient Transfer Learning
conference: Make sure to enter the correct conference title from your rights confirmation emai; 28 Oct–1 Nov, 2024; Melbourne, AUjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3680943isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer graphicsccs: Computing methodologies Image compression
@@ 1. Introduction
The importance of video compression techniques is driven by the need to efficiently store and transmit an ever-growing variety of video content. However, conventional video compression standards (e.g., H.266/VVC (Bross et al., 2021)) rely on manually designed modules that may no longer suffice for contemporary requirements. In recent times, Neural Video Compression (NVC) techniques (Li et al., 2023) that are optimized end-to-end with advanced neural modules have reached performance levels comparable to those of traditional codecs. Zhenghao Chen and Dong Xu are corresponding authors.
Email: zhenghao.chen@newcastle.edu.au, dongxu@hku.hk
@@
Despite achieving significant success, there are still some limitations to NVC in practical coding applications. The primary challenge is generalization, as most existing NVC models struggle to perform adaptively across different coding contents. This issue arises because most pre-trained NVC models are often optimized for a specific expected rate-distortion (RD) cost within a certain domain (e.g., online video resource (Xue et al., 2019)), leading to a notable domain gap when applied to cases outside the training domain (e.g., medical volumetric images). To mitigate the issue of such domain gaps, several content-adaptive methodologies (Pan et al., 2022; Yang et al., 2020a; Campos et al., 2019; Xu et al., 2023; Zhao et al., 2021; Gao et al., 2022; Van Rozendaal et al., 2021; Jourairi et al., 2022; Lam et al., 2020; Zou et al., 2021; Tsubota et al., 2023; Shen et al., 2023; Wang et al., 2022) have been introduced in the field of neural image compression (NIC). Many approaches involve updating the parameters of the encoder neural modules (i.e., online-updating strategy) to generate quantized compressed features with improved rate-distortion values. Such enhancement significantly boosts compression performance across various testing scenarios.
@@
Conversely, adopting the direct migration of update strategies from NIC to NVC frameworks still seems impractical. A primary obstacle lies in the evolution of NVC architectures, which now incorporate extensively complex modules (Li et al., 2021, 2022, 2023; Chen et al., 2023) with millions of parameters on the encoding side. Some early studies (Lu et al., 2020) have explored online-updating all parameters on the encoding side mirroring strategies used in content-adaptive NICs. Though it initially yielded performance gains in early light-weight NVC frameworks (Lu et al., 2019), applying such updates to the intricate architectures of contemporary NVCs will result in immense computational demands for encoding. Recent investigations (Tang et al., 2023) have also explored updating solely the encoding components of motion to circumvent the necessity of optimizing all encoding-side parameters. However, given that motion usually represents just a minor portion of the total bit-rates, confining updates to merely motion components substantially limits the scope for performance enhancements.
@@
Another challenge limiting the effectiveness of content-adaptive NVC is error accumulation. Video compression heavily relies on temporal redundancy. i.e., the compression performance of the current frame is greatly dependent on the reconstruction quality of the previously compressed frame. However, the online updating strategy, which aims to optimize the rate-distortion value, could simultaneously minimize both bit-rate and reconstruction quality for current coding frames. This might negatively impact subsequent coding frames through increased error accumulation. Such a dynamic underlines the difficulty of achieving both high compression performance and maintaining quality across frames.
@@
To tackle the aforementioned challenges, we propose a novel content-adaptive NVC method, termed Group-aware Parameter-efficient Updating (GPU). First, leveraging the breakthroughs in Parameter-Efficient Transfer Learning (PETL) (Ding et al., 2023; He et al., 2021; Houlsby et al., 2019) across various vision and language tasks, which demonstrates the efficiency of adapting a pre-trained model to new domains by fine-tuning a minimal set of parameters, our approach incorporates a low-rank adaptation style (Hu et al., 2021b) module, namely encoding-side adapters. These adapters are ingeniously crafted to stimulate specific components within the comprehensive encoding modules. By focusing optimization efforts on these adapters, which constitute a significantly smaller portion of the encoding modules’ parameters, we can then streamline optimization and reduce costs dramatically. Moreover, we delve into innovative ways to integrate these adapters into the NVC framework, exploring both serial and parallel configurations. The serial adapter is strategically placed immediately after a module, enhancing its processing, while the parallel adapter is woven into the module through a skip-connection layout, offering a blend of original and adapted processing pathways. Combining such strategic placements of adapters allows for the modification of less than 8% of the parameters in our encoding module, markedly boosting compression efficiency with minimal modification.
@@
Second, to mitigate error accumulation, we introduce a group-aware optimization strategy. This method diverges from previous methods (Abdoli et al., 2023; Lu et al., 2020; Tang et al., 2023; Van Rozendaal et al., 2021), which concentrated on updating a narrow sequence of consecutive frames. Instead, our approach considers the entire Group of Pictures (GoP) during the updating process. Specifically, when compressing an individual frame, we optimize the entire GoP, which it belongs, to minimize the errors in individual reconstructed frames. Furthermore, to circumvent the significant computational demand that group-based updating could impose on the processing of high-resolution videos, we implement a patch-based GoP updating strategy. It spatially segments each GoP into multiple patch-based GoPs, which will be updated sequentially. In general, this approach ensures that the updates made to this framework do not adversely impact the predictive coding process of other frames within the GoP, thus maintaining the integrity and efficiency of the compression across the entire group.
@@
By integrating two above strategies, we incorporate our GPU into the latest NVC framework (Li et al., 2023), significantly improving its coding performance across six most representative standard video compression benchmarks, UVG (UVG, [n. d.]), MCL-JCV (Wang et al., 2016) and HEVC (Sullivan et al., 2012) Class B, C, D and E datasets. Moreover, to validate the adaptability of our content-adaptive method in various contents, we also test it on compressing a medical volumetric image benchmark, ACDC (Bernard et al., 2018), which serves the most important cardiac diagnostics dataset with sequential MRI slices. Our framework not only outperforms traditional video codecs, such as H.266/VVC (Bross et al., 2021), but also surpasses previous content-adaptive NVC methods (Lu et al., 2020; Tang et al., 2023) in performance on video benchmarks and adaptability on medical benchmark. Our contributions can be summarized as follows:
@@
  • Implementation of a group-aware optimization strategy to mitigate error accumulation by considering an entire GoP during online updating, ensuring consistent and improved predictive coding performance across frames.
  • Integration of our approach to enhance the coding performance of the latest NVC methods, demonstrating superior results over both traditional video codecs and existing content-adaptive NVC methods across six key video and one medical volumetric image compression benchmarks.
@@ 2.1. Neural Video Compression
Current methods in neural video compression can be intuitively categorized into two main types: deep residual coding and deep contextual coding. The first-category approach is inspired by established video compression standards (Sullivan et al., 2012; Bross et al., 2021), which involves predictive coding (e.g., motion compensation) and the encoding of residual data. A notable pioneer in this category is DVC (Lu et al., 2019), which utilizes Convolutional Neural Networks (CNNs) throughout the traditional residual coding process. Following this, numerous works (Hu et al., 2020, 2021a; Chen et al., 2022b; Lin et al., 2020; Yang et al., 2020b; Agustsson et al., 2020) have enhanced this framework by incorporating more robust modules and sophisticated approaches. For instance, FVC (Hu et al., 2021a) implements all coding operations in the feature space to improve compression efficiency.
@@
In the deep contextual coding category (Lombardo et al., 2019; Habibian et al., 2019; Li et al., 2021; Sheng et al., 2022; Li et al., 2022, 2023, 2024; Chen et al., 2023, 2022a; Mentzer et al., 2022), research extends generative-based neural image compression (NIC) techniques by developing spatio-temporal conditional entropy models that leverage both spatial and temporal contexts. In the early studies, some works (Lombardo et al., 2019; Habibian et al., 2019) have introduced spatial-temporal conditional models by maintaining a latent temporal variable produced from all previous compressed frames. Unlike earlier approaches that rely on accumulating all available information, DCVC family (Li et al., 2021; Sheng et al., 2022; Li et al., 2022, 2023, 2024) introduces a series of methods that employ motion compensation from the consecutive compressed frame to derive temporal context directly, which has been further refined by progressive introducing new methods enhance such as hybrid entropy model (Li et al., 2022). The latest iteration of DCVC (i.e., DCVC_DC (Li et al., 2023)) has shown to surpass the compression performance of the conventional video coding standard, H.266/VVC (Bross et al., 2021).
@@
Though above methods have achieved notable success, their performances are often constrained across diverse content. This limitation stems primarily from the fact that many NVC frameworks are pre-trained on specific video domains (Xue et al., 2019), which restricts their generalizability and adaptability to a wide range of content. To address this challenge, this work introduces a content-adaptive NVC method named GPU, which is seamlessly integrated into the latest NVC framework, DCVC_DC. This integration significantly enhances its compression performance, demonstrating the potential of content-adaptive approaches in overcoming the generalizability limitations of existing NVC frameworks.
2.2. Content-Adaptive Neural Compression
@@
Content-adaptive methods play essential roles in both NIC (Pan et al., 2022; Yang et al., 2020a; Campos et al., 2019; Xu et al., 2023; Zhao et al., 2021; Gao et al., 2022; Van Rozendaal et al., 2021; Jourairi et al., 2022; Lam et al., 2020; Zou et al., 2021; Tsubota et al., 2023; Chen et al., 2021; Shen et al., 2023; Wang et al., 2022; Han et al., 2024; Liu et al., 2023, 2024; Lv et al., 2023) and NVC (Lu et al., 2020; Tang et al., 2023; Abdoli et al., 2023; Van Rozendaal et al., 2021), which is to refine the latent representation or coding parameters during the encoding process. In this way, the model can adapt to current coding samples and reduce the domain shift between training and testing domains, which enhances the adaptability when compressing various types of testing data.
@@
Those approaches have yielded significant accomplishments in NIC. For instance, Campos et al. (Campos et al., 2019) refined latent variables through back-propagation, a technique that was further enhanced by Yang et al. (Yang et al., 2020a) through the incorporation of a differentiable quantization module. Concurrently, some researchers have pursued decoder-based strategies; for example, Shen et al. (Shen et al., 2023) performed decoder updating using adapters. Nonetheless, it’s important to note such decoder-side adaptation strategies typically use additional bit-rates for storing and transmitting decoder-side parameters.
@@
In the field of NVC, the development of content-adaptive strategies is still in its infancy. Lu et al. (Lu et al., 2020) introduced the first method of this kind by updating all parameters on the encoding side, drawing parallels with most content-adaptive approaches in NIC. This method was seamlessly integrated into early NVC framework (Lu et al., 2019), which typically featured light-weight architectures. However, as more advanced modules (e.g., Transformers (Vaswani et al., 2017)) are incorporated to enhance the performance, modern NVC frameworks (Li et al., 2021, 2022, 2023, 2024; Chen et al., 2023; Hu et al., 2022) have become significantly more complex. Updating the entire encoder is now impractical due to the substantial computational load and increased encoding time. To address this issue, Tang et al. (Tang et al., 2023) developed a novel methodology for online and offline updating, specifically targeting the compression of the motion. This approach requires motion information produced by H.266/VVC (Bross et al., 2021) as the auxiliary ground-truth. Yet, since motion data represents only a small fraction of the overall bit-rates, optimizing motion alone is insufficient for significant performance gains.
@@
In contrast, our method employs a group-aware and parameter-efficient updating strategy to overcome the aforementioned challenges. First, we jointly optimize the entire GoP when producing the bitstream for a single frame, which ensures that updating a particular frame does not lead to significant error accumulation. Second, we implement an encoder-side adaptor, which allows us to avoid optimizing the entire encoder network with those comprehensive modules. This adaptation makes the updating process more efficient in terms of parameter use.